### 05 Dec 2023 v1

@@ Abstract
In the realm of large language models (LLMs), enhancing instruction-following capability often involves curating expansive training data. This is achieved through two primary schemes: i) Scaling-Inputs: Amplifying (input, output) pairs per task instruction, aiming for better instruction adherence. ii) Scaling Input-Free Tasks: Enlarging tasks, each composed of an (instruction, output) pair (without requiring a separate input anymore). However, LLMs under Scaling-Inputs tend to be overly sensitive to inputs, leading to misinterpretation or non-compliance with instructions. Conversely, Scaling Input-Free Tasks demands a substantial number of tasks but is less effective in instruction-following when dealing with instances in Scaling-Inputs. This work introduces Muffin, a new scheme of instruction-following dataset curation. Specifically, we automatically Scale Tasks per Input by diversifying these tasks with various input facets. Experimental results across four zero-shot benchmarks, spanning both Scaling-Inputs and Scaling Input-Free Tasks schemes, reveal that LLMs, at various scales, trained on Muffin generally demonstrate superior instruction-following capabilities compared to those trained on the two aforementioned schemes. [^1] All the code and data are available at our project page: https://renzelou.github.io/Muffin/
@@ 5 Experimental Setup
Implementation Details. Following previous works (Wang et al., 2023b; Honovich et al., 2022; Yin et al., 2023), we finetune T5-LM (Raffel et al., 2020) of different scales (3B and 11B) on our Muffin. We provide more implementation details (e.g., hyper-parameters) in Appendix G.
@@ 6.1 Automatic Evaluation
Table 2 shows the main results of finetuning T5-3B and T5-11B on different datasets.
@@
Indirect Comparison. When considering the comparison with SuperNI, our Muffin can still get a comparable or even better performance under some metrics, across different model sizes. Considering the SuperNI is crafted by humans, these results demonstrate the highly promising instruction-following capacity of the models tuned on Muffin. Notably, Muffin constantly achieves better performances on the generation tasks of SuperNI-Test than SuperNI-Train (and T0 and T0++) while maintaining a comparable classification performance, indicating the efficiency of our instruction brainstorming and classification expansion strategies.
@@ 6.2 Human Evaluation
Table 3 shows the results. Compared with the strongest results among the 8 baselines, Muffin mostly results in better human acceptance ratios with large margins (except for the tiny lower result on T0-Eval). According to our further analysis, we found that the other baselines often misunderstand the task objective, leading to meaningless outputs (e.g., copying some pieces of input text, or violating the format required by the instruction). While Muffin is more likely to produce useful outputs to solve the task. Overall, Muffin yields an substantial improvement on the average acceptance ratio of four benchmarks. We provide some representative cases of different systems’ outputs in Table 13.
@@ 6.3 Analyses
In addition to the automatic & human evaluations in the above two subsections, we further conduct some analyses to answer the remaining questions: ($\mathcal{Q}_{1}$) How much does each technical module (i.e., instruction brainstorm, rematching, classification expansion) contribute to the overall performance? ($\mathcal{Q}_{2}$) Does our method’s superior performance stem from latent task similarities or even leaking towards the evaluation tasks? ($\mathcal{Q}_{3}$) Does the data size influence the conclusion?
@@ Methods
Figure 5: The scaling trends comparison between SuperNI and our Muffin.
@@
It is possible for LLMs to generate the tasks leaking the evaluation benchmarks, violating the zero-shot setting. Therefore, we randomly sample 200 task instructions from each benchmark and use Sentence Transformers (Reimers & Gurevych, 2019) [^7] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2/tree/main to convert all the training instructions and the sampled test instructions into embeddings. We report the average cosine similarity between each training dataset and evaluation benchmark in Figure 5. First, we can find that all the Scaling-Inputs datasets, namely SuperNI, Dynosaur and Unnatural Instruct, show relatively high instruction semantic similarities with SuperNI-Test (Scaling-Inputs as well). Meanwhile, though there are rematched instructions sourced from SuperNI in our Muffin, Muffin still maintain a relatively low similarity with SuperNI-Test, owing to the diverse task distribution as discussed in § 4. In a word, our Muffin demonstrates relatively low instruction similarities across all four benchmarks, proving the validity of our previous zero-shot instruction-following superiority.
@@ 7 Conclusion and Broader Impact
  • Appendix J: Data Cases with Human-Evaluation Validity
  • Appendix K: System Output Cases on the Evaluation Benchmarks
@@ Appendix D Data Statistics
We report the length distribution of the inputs, instructions, and outputs of Muffin in Figure 6. All the length is counted by words (we follow previous works (Ma et al., 2021b; Wang et al., 2023b) using NLTK package [^9] https://pypi.org/project/nltk/ to conduct the word tokenization). Our dataset covers a diverse length distribution (see the maximum and minimum length of each figure), however, the most common length is still highly focused in a certain range, as the average length reported in Table 1.
@@ Appendix G Details of Tuning Muffin
All of our implementations are based on the Tk-Instruct [^11] https://github.com/yizhongw/Tk-Instruct that uses HuggingFace transformers (Wolf et al., 2019) [^12] https://pypi.org/project/transformers/ to implement all the models. Besides, all the models’ weights are downloaded from HuggingFace repositories. [^13] We use T5-3B from https://huggingface.co/t5-3b/tree/main and T5-11B from https://huggingface.co/t5-11b/tree/main. We fine-tune all the models on Muffin with 2 epochs. When fine-tuning T5-3B, we set the learning rate as $55e-$ 5 with batch size 6. As for T5-11B, we set the learning rate as $51e-$ 5 with batch size 1 and 12 gradient accumulation steps. We truncate the inputs to 1024 tokens and limit the output length to 128 tokens. All the above hyper-parameters are tuned on the validation set of SuperNI, and all the experiments are done on NVIDIA A100 with 80 GPU memories. We utilize DeepSpeed ZeRO stage 2 and 3 for single-GPU and multiple-GPUs tuning, respectively. We will release our training code with detailed document for future research.
@@ Appendix H Details of Baselines
We elaborate on more details of each baseline dataset used in our experiments. Following Honovich et al. (2022), all the hyper-parameters of fine-tuning baselines are trialed on the validation set of SuperNI. We truncate all the input length into 1024 and the output length into 128 (except those datasets designed for long response generation), which is the same as fine-tuning our Muffin. For a fair comparison, analogous to (Yin et al., 2023), we keep similar data size for all systems (60k $\sim$ 80k) with downsampling when necessary, and report the average performance of three seeds. When evaluating the models, we use the official scripts (Rajpurkar et al., 2016)  [^14] https://github.com/yizhongw/Tk-Instruct/blob/main/src/compute_metrics.py to calculate the metrics, which are widely adopted by the previous works (Wang et al., 2021; Lou et al., 2021).
@@
Scaling-Inputs is human-crafted high-quality instruction dataset, covering more than 1,600 tasks either from existing NLP benchmarks or created by human expertise. [^23] https://instructions.apps.allenai.org/ When training on SuperNI, we follow the official implementation tunning models on the 756 training tasks with a sequential multi-task learning scheme (Ma et al., 2021a; Zhang et al., 2023b). We use the official code and hyper-parameter settings of (Wang et al., 2022) to reproduce the performances.
@@ Appendix I Training and Evaluation Prompts
Appendix J Data Cases with Human-Evaluation Validity
@@ 2. Add crushed ice and blend again until smooth […]
Appendix K System Output Cases on the Evaluation Benchmarks

### 15 Mar 2024 v3

@@ Abstract
In the realm of large language models (LLMs), enhancing instruction-following capability often involves curating expansive training data. This is achieved through two primary schemes: i) Scaling-Inputs: Amplifying (input, output) pairs per task instruction, aiming for better instruction adherence. ii) Scaling Input-Free Tasks: Enlarging tasks, each composed of an (instruction, output) pair without requiring a separate input anymore. However, LLMs under Scaling-Inputs tend to be overly sensitive to inputs, leading to misinterpretation or non-compliance with instructions. Additionally, Scaling Input-Free Tasks demands a substantial number of tasks but is less effective in instruction-following when dealing with instances in Scaling-Inputs. This work introduces Muffin, a new scheme of instruction-following dataset curation. Specifically, we automatically Scale Tasks per Input by diversifying these tasks with various input facets. Experimental results across four zero-shot benchmarks, spanning both Scaling-Inputs and Scaling Input-Free Tasks schemes, reveal that LLMs, at various scales, trained on Muffin generally demonstrate superior instruction-following capabilities compared to those trained on the two aforementioned schemes. [^1] All the code and data are available at our project page: https://renzelou.github.io/Muffin/
@@ 5 Experimental Setup
Implementation Details. Following previous works (Wang et al., 2023b; Honovich et al., 2022; Yin et al., 2023), we finetune models with various architectures on Muffin, including encoder-decoder T5-LM (Raffel et al., 2020) and decoder-only Llama (Touvron et al., 2023a). We provide more implementation details (e.g., hyper-parameters) in Appendix G.
@@ 6.1 Automatic Evaluation
Table 2 shows the main results of fine-tuning T5-3B and T5-11B on different datasets. Additionally, we add the results based on Llama2 in Table 3 to further enhance our conclusion.
@@
[…]
SuperNI-Test MMLU T0-Eval BBH
Models Data Size
EM
ROUGE-L
ROUGE-L
EM EM EM
Human Annotated Data (indirect comparison)
[…]
Generated Data (direct comprison)
Dolly 15k 9.96 43.58 27.25 0.39 22.29 7.76
LongForm 23k 4.30 41.30 19.07 0.12 0.72 5.27
Alpaca 52k 33.34 51.67 43.65 36.01 40.39 21.72
Alpaca-GPT4 52k 18.27 44.27 33.50 1.01 6.29 2.20
WizardLM 68k 10.52 43.36 27.27 0.29 7.20 4.24
[…]
Unnatural Inst. 68k 37.63 50.23 46.03 6.69 8.35 5.05
Dynosaur 66k 44.35 49.34 47.08 17.26 34.59 7.11
[…]
@@
Indirect Comparison. When considering the comparison with SuperNI, our Muffin can still get a comparable or even better performance under some metrics, across different models and model sizes. Considering the SuperNI is crafted by humans, these results demonstrate the highly promising instruction-following capacity of the models tuned on Muffin. Notably, Muffin constantly achieves better performances on the generation tasks of SuperNI-Test than SuperNI-Train (and T0 and T0++) while maintaining a comparable classification performance, indicating the efficiency of our instruction brainstorming and classification expansion strategies.
@@ 6.2 Human Evaluation
Table 4 shows the results. Compared with the strongest results among the 8 baselines, Muffin mostly results in better human acceptance ratios with large margins (except for the tiny lower result on T0-Eval). According to our further analysis, we find that the other baselines often misunderstand the task objective, leading to meaningless outputs (e.g., copying some pieces of input text, or violating the format required by the instruction). While Muffin is more likely to produce useful outputs to solve the task. Overall, Muffin yields a substantial improvement on the average acceptance ratio of four benchmarks. We provide some representative cases of different systems’ outputs in Table 18.
@@ 6.3 Analyses
In addition to the automatic & human evaluations in the above two subsections, we further conduct some analyses to answer the remaining questions: ($\mathcal{Q}_{1}$) How much does each technical module (i.e., instruction brainstorm, rematching, classification expansion) contribute to the overall performance? ($\mathcal{Q}_{2}$) Does our method’s superior performance stem from latent task similarities or even leaking towards the evaluation tasks? ($\mathcal{Q}_{3}$) Does the data size influence the conclusion? ($\mathcal{Q}_{4}$) Does mixing Muffin with human-crafted instruction instances boost the performance?
@@ Methods
It is possible for LLMs to generate the tasks leaking the evaluation benchmarks, violating the zero-shot setting. Therefore, we randomly sample 200 task instructions from each benchmark and use Sentence Transformers (Reimers & Gurevych, 2019) [^7] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2/tree/main to convert all the training instructions and the sampled test instructions into embeddings. We report the average cosine similarity between each training dataset and evaluation benchmark in Figure 9. We also follow the previous works  (Yin et al., 2023) using ChatGPT to estimate the task overlap in Table Renze. First, we can find that all the Scaling-Inputs datasets, namely SuperNI, Dynosaur and Unnatural Instruct, show relatively high instruction semantic similarities with SuperNI-Test (Scaling-Inputs as well). Meanwhile, though there are rematched instructions sourced from SuperNI in our Muffin, Muffin still maintain a relatively low similarity with SuperNI-Test, owing to the diverse task distribution as discussed in § 4. In a word, our Muffin demonstrates relatively low instruction similarities across all four benchmarks, proving the validity of our previous zero-shot instruction-following superiority.
Figure 4: The scaling trends comparison between Muffin and the previous baseline datasets (average performances on all four benchmarks).
@@
Regarding the cost efficiency of the previous works, it’s usually convenient to collect extensive instruction-following instances with previous paradigms (Yin et al., 2023; Wu et al., 2023). Therefore, is it worthwhile to try the proposed paradigm rather than simply gathering more instances from the other paradigms? To answer this question, for all those baseline datasets from the “direct comparison” in Table 2, we randomly sample subsets from them and train T5-3B on the subsets to show the performance trends (10%, 30%, 50%, 80%, 100%). As shown in Figure 4, Muffin exceeds the baselines by a noteworthy margin (average scores on four evaluation benchmarks). Other baselines may only be comparable to our data results when they continue to be scaled to several times the size of our data. More importantly, the performances of some datasets even decrease after scaling to a larger size (perhaps due to the noise in these LLM-synthetic datasets), such as Self-Instruct and Dynosaur. Therefore, we conjecture that our paradigm is more efficient than simply collecting more data from the other paradigms. We also conduct the scaling comparison with SuperNI in Appendix J.1.
Figure 5: The performance of the mixture of Muffin and SuperNI.
Mixing Muffin with SuperNI to answer $\mathcal{Q}_{4}$.
As demonstrated by the previous works, it’s possible to further improve the generalization performance after mixing the LLM-synthetic and human-annotated instruction instances (Yin et al., 2023). Thus, we mix Muffin with the training set of SuperNI, w.r.t. various proportions, where the proportion means how many instances are from Muffin. Then, we train T5-3B on these mixtures and report the average performances on the four benchmarks. As illustrated in Figure 5, interestingly, after mixing Muffin and the human-annotated SuperNI, the performance drops even worse than only using Muffin for training. We conjecture possible reasons for this phenomenon: the dataset paradigm does affect the learning efficiency — different dataset paradigms do obviously have various impacts on the model’s performance. Therefore, combining two paradigms can even hurt the model’s instruction-following performance, meaning the paradigm is critical. Meanwhile, the effect of dataset paradigms is even greater than that of data quality (i.e., adding a small proportion of high-quality data from other paradigms even harms Muffin’s performance), further implying the proposed paradigm’s effectiveness.
@@ 7 Conclusion and Broader Impact
  • Appendix J: Further Analyses
@@ Appendix D Data Statistics
We report the length distribution of the inputs, instructions, and outputs of Muffin in Figure 6. All the length is counted by words (we follow previous works using NLTK package [^9] https://pypi.org/project/nltk/ to conduct the word tokenization). Our dataset covers a diverse length distribution (see the maximum and minimum length of each figure), however, the most common length is still highly focused in a certain range, as the average length reported in Table 1.
@@ Appendix G Details of Tuning Muffin
All of our implementations are based on HuggingFace transformers (Wolf et al., 2019). [^11] https://pypi.org/project/transformers/ For the implementation of tuning T5 models, we adopt the source training code of SuperNI; [^12] https://github.com/yizhongw/Tk-Instruct for fine-tuning the Llama2, we use the open-sourced Alpaca-Lora, [^13] https://github.com/tloen/alpaca-lora where we apply LoRA (Hu et al., 2022) to relieve the high computational cost. Besides, all the models’ weights are downloaded from HuggingFace repositories. [^14] We use T5-3B from https://huggingface.co/t5-3b/tree/main, T5-11B from https://huggingface.co/t5-11b/tree/main, and Llama2-13B from https://huggingface.co/meta-llama. We fine-tune T5 on Muffin with 2 epochs. When fine-tuning T5-3B, we set the learning rate as $55e-$ 5 with batch size 6. As for T5-11B, we set the learning rate as $51e-$ 5 with batch size 1 and 12 gradient accumulation steps. All the above hyper-parameters are tuned on the validation set of SuperNI. While we fine-tune Llama2 on all the datasets 3 epochs with batch size 18, and we set learning rate $=1e-4$, $lora_{r}=8$, $lora_{\text{alpha}}=16$. Since the generation API provided by HugggingFace cannot support efficient batched evaluation, we fix the evaluation batch size to 1 for all the datasets. We truncate the inputs to 1024 tokens and limit the output length to 128 tokens, with beam search size $=1$ (greedy decoding).
All the experiments are done on NVIDIA A100 with 80 GPU memories. When fine-tuning T5 models, we utilize DeepSpeed ZeRO stage 2 and 3 for single-GPU and multiple-GPU tuning, respectively.
@@ Appendix H Details of Baselines
We elaborate on more details of each baseline dataset used in our experiments. Following Honovich et al. (2022), all the hyper-parameters of fine-tuning baselines are trialed on the validation set of SuperNI. We truncate all the input length into 1024 and the output length into 128 (except those datasets designed for long response generation), which is the same as fine-tuning our Muffin. For a fair comparison, analogous to (Yin et al., 2023), we keep similar data size for all systems (60k $\sim$ 80k) with downsampling when necessary, and report the average performance of three seeds. When evaluating the models, we use the official scripts (Rajpurkar et al., 2016)  [^15] https://github.com/yizhongw/Tk-Instruct/blob/main/src/compute_metrics.py to calculate the metrics, which are widely adopted by the previous works.
@@
Scaling-Inputs is human-crafted high-quality instruction dataset, covering more than 1,600 tasks either from existing NLP benchmarks or created by human expertise. [^24] https://instructions.apps.allenai.org/ When training on SuperNI, we follow the official implementation tunning models on the 756 training tasks with a sequential multi-task learning scheme. We use the official code and hyper-parameter settings of (Wang et al., 2022) to reproduce the performances.
@@ Appendix I Training and Evaluation Prompts
Appendix J Further Analyses
Figure 9: The scaling trends comparison between SuperNI and our Muffin. The performances are based on T5-3B.
J.1 Scaling Trends Comparison with SuperNI
Table 13: The task category overlap between different training datasets and evaluation benchmarks, estimated by ChatGPT.
SuperNI-Test MMLU T0-Eval BBH
Self-Inst. 11.62% 10.37% 11.27% 5.29%
Unnatural Inst. 22.27% 16.91% 18.39% 4.76%
Dynosaur 12.08% 13.49% 15.20% 1.27%
Muffin 16.01% 13.05% 8.63% 2.53%
J.2 The Task Overlap Estimation
Beyond the task similarity analysis provided in § 6.3, we follow previous work prompting ChatGPT to further quantify the task category overlap between the evaluation benchmarks and different training datasets (Yin et al., 2023). Given an instance from a training dataset, we ask ChatGPT whether it belongs to any task categories from the evaluation benchmark. For example, the test set of SuperNI has 12 task categories, and we request ChatGPT to perform a 13-way classification (one “None of above” label) on the training instances. Finally, we report how many instances from this training dataset contain task leakage.
Table 13 illustrates the results. Here, we only report the overlap results of Muffin and those competitive baseline datasets according to Table 2. Though there are some shifts in the leakage ranking compared with the semantic-based analysis in § 6.3 (see Figure 9), the overall conclusion is still the same — Muffin demonstrates relatively low evaluation task leakage across all four benchmarks.
Table 14: Create Muffin with the input sourced from Dynosaur and Unnatural Instruct (namely Muffin-Dynosaur and Muffin-Unnatural). All the datasets here have around 16k instances. The experiment is based on T5-3B.
Models SupnerNI-Test (overall) MMLU (ACC) MMLU (EM) T0-Eval (ACC) T0-Eval (EM) BBH (EM)
Unnatural Inst. 37.12 24.35 22.33 45.93 35.88 8.69
Dynosaur 32.55 26.88 25.26 38.56 39.13 12.09
Muffin-Unnatural 38.86 33.12 24.73 46.08 43.63 12.91
Muffin-Dynosaur 38.23 33.34 22.59 45.3 44.01 13.53
Muffin 40.45 33.24 25.39 42.68 40.99 13.44
J.3 Using the Inputs Sourced from the Other LLM-synthetic Datasets
As introduced in § 3, the creation of Muffin utilizes the human-crafted inputs from SuperNI (i.e., “instruction brainstorm”). While those baselines with Scaling-Inputs paradigm either utilize inputs from other human-crafted datasets (e.g., Dynosaur) or let LLMs to synthesize inputs. To investigate whether the proposed “instruction brainstorm” method can be extended to other input sources, we gather the inputs from other LLM-generated datasets (including Dynosaur and Unnatural Instruct), and apply our “insruction brainstorm” method to them. Finally, we create two new small-scale datasets (around 16k instances), namely Muffin-Dynosaur and Muffin-Unnatural, and compare their performances with Dynosaur and Unnatural Instruct (in the same size). Table 14 shows the results, where we observe that Muffin consistently demonstrates superior generalization performance, regardless of the text sources employed. It further suggests that Muffin’s robustness to follow instructions is not reliant on the input resources but rather stems more from our crucial contribution—the diversified instructions per input. Therefore, we anticipate that our method can be ideally applied to any input resources while still crafting diverse task instructions.
J.4 The Explanability of the Generated Facets
As mentioned in § 3, the proposed “instruction brainstorm” method drives LLMs to brainstorm novel instructions based on the synthetic facets. To this end, it’s unclear whether the synthetic facets are explainable and reliable for the subsequent instruction generation due to the LLM’s hallucinations (Xie et al., 2023; Zhang et al., 2023b). Therefore, in this section we conduct several additional analysis on the generated facets in Muffin, including statistics, quality, and diversity of the facets.
Table 15: The statistics of generated facets.
facets statistics
# of inputs 1,463
# of instructions (from facets-based brainstorm) 33,720
# of facets 11,382
avg # of facets per input 7.78
max # of facets per input 30
min # of facets per input 2
avg # of instructions per facet 2.96
max # of instructions per facet 18
min # of instructions per facet 1
Facets Statistics.
Beyond the statistics of instructions in Table 1, we report additional statistics related to facets in Table 15 (the statistics after conducting “instruction filtering”).
Factes Quality.
To estimate the quality and reliability of the generated facets, we conduct a further human verification on the facets. Specifically, we randomly collect 100 inputs from Muffin, along with their corresponding facets (889 facets in total) and instructions (2,265 instructions in total). Then, we ask an experienced human annotator to evaluate the following two metrics:
  • Input-to-facet correctness: whether each facet correctly described the given input (reflecting the quality and reliability of facts).
  • Facet-to-instruction correctness: whether the subsequent instructions are reasonably related to the given facet (reflecting the utility of facets).
The final results are shown in Table 16. These results demonstrate the good quality and utility of the facets generated by our method. It’s also worth noting that, after we conducted ”instruction filtering” (as introduced on § 3), the input-to-facets correctness increased from 83.28% to 90.78%, proving the effectiveness of the naive filtering method that can jointly filter a considerable amount of hallucinated facets.
Table 16: The quality evaluation results of generated facets.
input-to-facet correctness 90.78%
facet-to-instruction correctness 85.22%
Facets Diversity.
As for the diversity of the generated facets, we ask the human annotator to discover the topic categories from the facets (i.e., clustering instead of classification), because it’s hard to predefine the categories that may introduce personal bias. We used the same instances as in the previous analysis (100 inputs and 889 facets). The human annotator is asked to summarize a “keyword” (topic category) for each LLM-generated facet. If the current category has already appeared before, the annotator also has to “cluster” those facets that belong to the same category together. In doing so, more and more novel categories can be discovered.
We then calculate the following two metrics:
  • Intra-diversity: how many unique categories there were for each input’s facets, and reported an averaged ratio among all the inputs.
  • Inter-diversity: overall categories distribution among all the 889 facets, namely the global frequency of each unique facet.
We find that the generated facets achieve 91.55% intra-diversity, meaning the facets of the same input are pretty diverse. Meanwhile, there are a total of 276 unique categories out of the 889 facets, which implies a high inter-diversity of the facets.
J.5 Data Cases with Human-Evaluation Validity
@@ 2. Add crushed ice and blend again until smooth […]
J.6 System Output Cases on the Evaluation Benchmarks